1.     The density of a substance is 2500 kg/m3. What volume will 250 kg of this substance occupy?

A.   0.001 m3                                                                                                     C. 0.10 m3

B.    0.01 m3                                                                    D. 1.0 m3

 

2.     In a typical hydraulic press, a force of 20 N is exerted on the small piston of area 0.050 m². The force exerted by the large piston on the load if it has an area of 0.50 m² will be

A.   200 N                                                                        C. 50 N

B.    100 N                                                                        D. 10 N 


1
Expert's answer
2022-02-07T15:27:15-0500

1.

"V=\frac{m}{\rho}=\frac{250\: \rm kg}{2500\: \rm kg/m^3}=0.1\:\rm m^3"

Answer: C

2.

"F_2=F_1\frac{A_2}{A_1}=\rm20\: N*\frac{0.50\: m^2}{0.050\: m^2}=200\: N"

Answer: A